# Asymgenesis

### v2

@@ Abstract
We present a framework based on the standard type-I seesaw model that relates the baryon asymmetry of the universe to the dark matter (DM) density. The framework, which we name Asymgenesis, relies on the presence of primordial charge asymmetries seeded either in the dark sector or in the visible sector. A higher-dimensional portal operator reshuffles this initial asymmetry into both sectors, eventually resulting in a nonzero $B-L$ asymmetry and an asymmetric DM component. Compared to conventional asymmetric-dark-matter (ADM) schemes, our framework imposes far milder requirements on the portal interaction. In particular, the portal interaction need not violate $B-L$, and the temperature scales of efficient $B-L$ violation and efficient charge-transfer interaction mediated by the portal operator can be separated. We develop the formalism in detail and argue that the flexibility of our framework enlarges the model-building landscape for ADM.
@@
Introduction —  Addressing the observed baryon asymmetry of the Universe (BAU) Planck:2018vyg ; ParticleDataGroup:2024cfk calls for physics beyond the Standard Model (SM). A leading paradigm for dynamically generating the BAU in the early universe is leptogenesis (LG) Fukugita:1986hr , which requires extending the SM field content by at least two right-handed neutrinos (RHNs). An attractive feature of LG is that RHNs can account for the observed nonzero masses of active neutrinos via the type-I seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Yanagida:1980xy ; Gell-Mann:1979vob ; Mohapatra:1979ia , thereby connecting the origin of the BAU to the origin of active neutrino masses Buchmuller:2005eh ; Bodeker:2020ghk . Most LG realizations rely on all three of Sakharov’s conditions Sakharov:1967dj for generating the BAU to emerge from the RHN sector Fukugita:1986hr ; Luty:1992un ; Covi:1996wh ; Pilaftsis:1997jf ; Akhmedov:1998qx ; Barbieri:1999ma ; Pilaftsis:2003gt ; Giudice:2003jh ; Buchmuller:2004nz ; Pilaftsis:2005rv ; Klaric:2020phc ; Klaric:2021cpi . A notable exception is wash-in leptogenesis (WILG), which does not require any source of charge-parity $(CP)$-violation from the RHN sector Domcke:2020quw ; Domcke:2022kfs ; Mojahed:2025vgf ; WILGRelated .
@@
The philosophy underlying WILG is to treat RHN interactions as spectator processes whose role is to provide a source of $B\!-\!L$ violation. In this way, RHNs are treated on par with electroweak sphalerons Kuzmin:1985mm , whose role in standard LG mechanisms is merely to serve as a source of $B+L$ violation. Meanwhile, $CP$ violation is attributed to novel $CP$-violating dynamics at high temperatures above the mass of the heaviest RHN $T\gg M_{N}$, in a process referred to as chargegenesis (CG), see Refs. Domcke:2022kfs ; Schmitz:2023pfy ; Mukaida:2024eqi ; Mojahed:2024yus for concrete realizations. The role of CG is to provide $CP$-violating initial conditions for WILG, and can be viewed as a UV completion for the effective framework of WILG. In particular, the SM entails eleven linearly independent conserved global charges at high temperatures that could be produced during CG Domcke:2020quw . During WILG, RHNs act on the nontrivial chemical background induced by CG in a $B-L$ violating manner and drive the thermal plasma to an attractor solution featuring nonzero $B-L$, even if $B-L$ was conserved throughout CG Domcke:2020quw .
@@
Within the framework presented in this paper, we demonstrate that the initial conditions required for successful WILG can also be used to address the generation of the DM in our universe Planck:2018vyg ; Cirelli:2024ssz . In particular, if DM has a particle nature and consists primarily of either dark particles or dark anti-particles, then a primordial charge asymmetry in the SM sector can induce suitable initial conditions for ADM. Similarly, a primordial charge asymmetry in the dark sector can also induce initial conditions suitable for WILG. This idea is closely related to the standard ADM paradigm Kaplan:1991ah ; Nussinov:1985xr ; Kaplan:2009ag ; Petraki:2013wwa ; Zurek:2013wia , where the generation of the BAU and the DM abundance are tightly connected. In many ADM models, either a $B-L$ asymmetry or a DM asymmetry is generated at high temperatures, and subsequently transferred to the other sector Shelton:2010ta ; Haba:2010bm ; Buckley:2010ui ; Blennow:2010qp ; Bernal:2016gfn ; Hall:2019ank ; Hall:2021zsk ; Barman:2021tgt ; Asadi:2025vli ; ADMSeesaw . A standard transfer mechanism is provided by a higher-dimensional operator of the form $\mathcal{O}_{B-L}\mathcal{O}_{D}/M^{n}$, where $\mathcal{O}_{B-L}$ violates $B-L$ and $\mathcal{O}_{D}$ violates a global DM charge Zurek:2013wia . At high temperatures, the interaction rates induced by this operator are efficient, and the asymmetry is freely shared between the two sectors. As the temperature in the early universe drops, the interaction decouples, and an asymmetry is separately frozen in the dark and visible sectors. Finally, as the temperature drops below the mass of the DM particle, particle-antiparticle annihilations erase the symmetric component of the DM density, resulting in a DM density determined by the particle-antiparticle asymmetry and prediction of a $\mathcal{O}$(GeV) DM mass Zurek:2013wia .
However, as will be shown in this work, the restrictions on the higher-dimensional interaction transferring asymmetry between the two sectors can be significantly relaxed in the context of the WILG paradigm. Rather than violating $B-L$, it is sufficient if the interaction takes the form
@@
where $\mathcal{O}_{C}^{\rm{(SM)}}$ violates one of the conserved charges $C$ used in WILG and $\tilde{\Lambda}$ is a heavy mass scale. In general, $\mathcal{O}_{C}^{\rm{(SM)}}$ can be any gauge-invariant operator, but for concreteness, we will consider the case where it is renormalizable, and reduces to a SM Yukawa-type operator. We will also consider the simplest realization possible for the dark operator $\mathcal{O}^{(D)}_{X}=X^{p}$, where $X$ is a complex scalar asymmetric DM candidate, and $\mathcal{O}_{X}^{(D)}$ violates ”$X$-number” charge by $p$ units. Throughout, we assume that $X$ belongs to a dark sector that allows for sufficiently efficient annihilation of the symmetric abundance of $X$ such that the asymmetric component provides the dominant contribution to the DM density, and that the charge-transfer interaction in Eq. (1) provides the dominant source of $X$-number violation.
The framework presented here, which we name Asymgenesis, can be regarded as a flexible and significant extension of the ADM paradigm. In contrast to standard ADM mechanisms, Asymgenesis allows for a separation between the scale of charge-transfer between the dark and the visible sectors and $B-L$ violation. Indeed, Asymgenesis includes the intriguing possibility that the charge-transfer mechanism need not violate $B-L$. Instead, it is sufficient that any of the (global) SM charges that are approximately conserved at high temperatures are violated, allowing for many more options for charge transfer.
@@
Warmup —  We initiate our analysis by examining the transfer of charge between the visible and the dark sector. The charge asymmetry $q_{i}$ of a SM or DM particle species $i$ is defined as, $q_{i}=n_{i}-n_{\bar{i}}=g_{i}\mu_{i}T^{2}/6$, where $n_{i}$ $(n_{\bar{i}})$ denotes the (anti)particle number density of particle $i$ with chemical potential $\mu_{i}$ and multiplicity $g_{i}$. At high temperatures, the interaction described in Eq. (1) may establish thermal equilibrium between the two sectors, relating their chemical potentials $\mu_{i}$ Harvey:1990qw . Further, all SM and DM chemical potentials can be expressed as linear combinations of a set of linearly independent conserved charges $q_{C}=\bar{\mu}_{C}T^{2}/6$, e.g. by employing the linear-algebra formalism developed in the appendix of Ref. Domcke:2020quw . Here, $\bar{\mu}_{C}$ is introduced for notional convenience: $\bar{\mu}_{C}$ and $q_{C}$ are not conjugate to each other in the usual thermodynamical sense, but both describe the conserved charge. In other words, $\bar{\mu}_{C}$ corresponds to $q_{C}$, written in units of a chemical potential. To illustrate the concept of Asymgenesis, we consider a specific example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$, with $L_{e}$, $H$, and $e_{R}$ representing the first-generation lepton doublet, the SM Higgs doublet, and the right-handed electron, respectively. If the higher-dimensional operator is in thermal equilibrium, the following equilibrium condition relating the SM and the dark sector holds,
@@
While the operator $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ violates right-handed electron number, the interaction Eq. (1) conserves a combination of right-handed electron number and dark number. Consequently, the conservation of the right-handed electron charge, $q_{e}\sim\mu_{e}/T$, at high temperatures Campbell:1992jd ; Cline:1993vv ; Cline:1993bd ; Bodeker:2019ajh is modified as follows
@@
where we used $g_{X}=2$, which we employ throughout this paper. As an example, if the charge-transfer interaction decouples at temperatures of a few 100100100 TeV, then the ADM chemical potential can be written as
@@
The ADM abundance would in this case be determined by two factors: (i) the primordial asymmetries in $q_{e}$ and $q_{X}$, and (ii) the values of the flavored $B-L$ asymmetries, $\Delta_{\alpha}$, at the time of charge-transfer decoupling. If the lightest RHN, $N_{1}$, has a mass around $M_{N}\sim 10^{2-3}$ TeV and decouples while the portal interaction is still efficient, then WILG in the strong wash-in regime would lead to

### v3

# arXiv:2507.10655v3[hep-ph] 02 Mar 2026

@@ Abstract
We present a framework based on the standard type-I seesaw model that relates the baryon asymmetry of the universe to the dark matter (DM) density. The framework, which we name Asymgenesis, relies on the presence of primordial charge asymmetries seeded either in the dark sector or in the visible sector. A higher-dimensional portal operator reshuffles this initial asymmetry into both sectors, eventually resulting in a nonzero BLB-L asymmetry and an asymmetric DM component. Compared to conventional asymmetric-dark-matter (ADM) schemes, our framework imposes far milder requirements on the portal interaction. In particular, the portal interaction need not violate BLB-L, and the temperature scales of efficient BLB-L violation and efficient charge-transfer interaction mediated by the portal operator can be separated. We develop the formalism in detail and argue that the flexibility of our framework enlarges the model-building landscape for ADM.
@@
Introduction —  Addressing the observed baryon asymmetry of the Universe (BAU) Aghanim and others (2020); Navas and others (2024) calls for physics beyond the Standard Model (SM). A leading paradigm for dynamically generating the BAU in the early universe is leptogenesis (LG) Fukugita and Yanagida (1986), which requires extending the SM field content by at least two right-handed neutrinos (RHNs). An attractive feature of LG is that RHNs can account for the observed nonzero masses of active neutrinos via the type-I seesaw mechanism Minkowski (1977); Yanagida (1979, 1980); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980), thereby connecting the origin of the BAU to the origin of active neutrino masses Buchmuller et al. (2005b); Bodeker and Buchmuller (2021). Most LG realizations rely on all three of Sakharov’s conditions Sakharov (1967) for generating the BAU to emerge from the RHN sector Fukugita and Yanagida (1986); Luty (1992); Covi et al. (1996); Pilaftsis (1997); Akhmedov et al. (1998); Barbieri et al. (2000); Pilaftsis and Underwood (2004); Giudice et al. (2004); Buchmuller et al. (2005a); Pilaftsis and Underwood (2005); Klarić et al. (2021b, a). A notable exception is wash-in leptogenesis (WILG), which does not require any source of charge-parity $(CP)$-violation from the RHN sector Domcke et al. (2021, 2023); Mojahed et al. (2025); See also () for related ideas. .
@@
The philosophy underlying WILG is to treat RHN interactions as spectator processes whose role is to provide a source of BLB\!-\!L violation. In this way, RHNs are treated on par with electroweak sphalerons Kuzmin et al. (1985), whose role in standard LG mechanisms is merely to serve as a source of $B+L$ violation. Meanwhile, $CP$ violation is attributed to novel $CP$-violating dynamics at high temperatures above the mass of the heaviest RHN $T\gg M_{N}$, in a process referred to as chargegenesis (CG), see Refs. Domcke et al. (2023); Schmitz and Xu (2024); Mukaida et al. (2024); Mojahed et al. (2024) for concrete realizations. The role of CG is to provide $CP$-violating initial conditions for WILG, and can be viewed as a UV completion for the effective framework of WILG. In particular, the SM entails eleven linearly independent conserved global charges at high temperatures that could be produced during CG Domcke et al. (2021). During WILG, RHNs act on the nontrivial chemical background induced by CG in a BLB-L violating manner and drive the thermal plasma to an attractor solution featuring nonzero BLB-L, even if BLB-L was conserved throughout CG Domcke et al. (2021).
@@
Within the framework presented in this paper, we demonstrate that the initial conditions required for successful WILG can also be used to address the generation of the DM in our universe Aghanim and others (2020); Cirelli et al. (2024). In particular, if DM has a particle nature and consists primarily of either dark particles or dark anti-particles, then a primordial charge asymmetry in the SM sector can induce suitable initial conditions for ADM. Similarly, a primordial charge asymmetry in the dark sector can also induce initial conditions suitable for WILG. This idea is closely related to the standard ADM paradigm Kaplan (1992); Nussinov (1985); Kaplan et al. (2009); Petraki and Volkas (2013); Zurek (2014), where the generation of the BAU and the DM abundance are tightly connected. In many ADM models, either a BLB-L asymmetry or a DM asymmetry is generated at high temperatures, and subsequently transferred to the other sector Shelton and Zurek (2010); Haba and Matsumoto (2011); Buckley and Randall (2011); Blennow et al. (2011); Bernal et al. (2016); Hall et al. (2020, 2022); Barman et al. (2021); Asadi et al. (2025). [^1] See e.g. also Mahapatra et al. (2025); Borah et al. (2024); Takahashi and Wada (2026) for some recent work on ADM in seesaw models. A standard transfer mechanism is provided by a higher-dimensional operator of the form $\mathcal{O}_{B-L}\mathcal{O}_{D}/\Lambda^{n}$, where $\mathcal{O}_{B-L}$ violates BLB-L and $\mathcal{O}_{D}$ violates a global DM charge Zurek (2014). At high temperatures, the interaction rates induced by this operator are efficient, and the asymmetry is freely shared between the two sectors. As the temperature in the early universe drops, the interaction decouples, and an asymmetry is separately frozen in the dark and visible sectors. This standard picture of charge transfer via higher-dimensional operators is illustrated on the left in Fig. 1. Finally, as the temperature drops below the mass of the DM particle, particle-antiparticle annihilations erase the symmetric component of the DM density, resulting in a DM density determined by the particle-antiparticle asymmetry and prediction of a $\mathcal{O}$(GeV) DM mass Zurek (2014).
@@
Figure 1: Left: Schematic of standard ADM models with charge transfer mediated by higher-dimensional operator(s) (adapted from Ref. Zurek (2014)). Right: A schematic of charge transfer in Asymgenesis.
However, as will be shown in this work, the restrictions on the higher-dimensional interaction transferring asymmetry between the two sectors can be significantly relaxed in the context of the WILG paradigm. Rather than violating BLB-L, it is sufficient if the interaction takes the form
$\mathcal{O}_{C}^{\rm{(SM)}}\mathcal{O}_{X}^{(D)}/\tilde{\Lambda}^{n}+\text{h.c.}\,$, (1)
where $\mathcal{O}_{C}^{\rm{(SM)}}$ violates at least one of the conserved charges $C$ used in WILG [^2] In many cases, $\mathcal{O}_{C}^{\rm{(SM)}}$ needs not even violate $C$, as explained in the Appendix. and $\tilde{\Lambda}$ is a heavy mass scale. In general, $\mathcal{O}_{C}^{\rm{(SM)}}$ needs not be renormalizable or gauge invariant or Lorentz invariant as long as invariance is restored for the combination $\mathcal{O}_{C}^{\rm{(SM)}}\mathcal{O}_{X}^{(D)}$. However, in this letter, we will consider the simplest case where it is a Lorentz scalar, renormalizable, and gauge invariant, and thereby reduces to a SM Yukawa-type operator. To the best of our knowledge, this possibility has not been considered in the literature so far. In Appendix A, we show that by relaxing these requirements one recovers a broader landscape of ADM constructions based on charge transfer via higher-dimensional operators, and we illustrate how the framework introduced in this letter can be applied to generalize established charge-transfer mechanisms. We will also consider the simplest realization possible for the dark operator $\mathcal{O}^{(D)}_{X}=X^{p}$, where $X$ is a complex scalar asymmetric DM candidate, and $\mathcal{O}_{X}^{(D)}$ violates ”$X$-number” charge by $p$ units. In scenarios with $p>2$, potentially dangerous $X-\bar{X}$ oscillations Buckley and Profumo (2012); Cirelli et al. (2012); Tulin et al. (2012) are autmatically avoided if $X$ is charged under a dark $\mathbb{Z}_{p}$ symmetry. Throughout, we assume that $X$ belongs to a dark sector that allows for sufficiently efficient annihilation of the symmetric abundance of $X$ such that the asymmetric component provides the dominant contribution to the DM density, and that the charge-transfer interaction in Eq. (1) provides the dominant source of $X$-number violation.
The framework introduced here, denoted Asymgenesis, represents a unification of the conventional ADM scenario — formulated via higher-dimensional operators — with the type-I seesaw mechanism, establishing a common framework for neutrino mass generation and asymmetry transfer between sectors. The role of RHNs in Asymgenesis is the same as in WILG, namely to act on a nontrivial chemical background that features charge asymmetries in a BLB-L violating manner. In contrast to standard ADM mechanisms, Asymgenesis allows for a separation between the scale of charge transfer between the dark and the visible sectors and BLB-L violation. Indeed, Asymgenesis includes the intriguing possibility that the charge-transfer mechanism need not violate BLB-L. This possibility was also realized in Higgsogenesis Servant and Tulin (2013) where a nonzero Higgs chemical potential – induced by either a primordial BLB-L asymmetry or a primordial dark charge – mediates asymmetries between visible and dark sectors. Meanwhile, in Asymgenesis a primordial abundance of any of the (global) SM charges that are approximately conserved at high temperatures suffices. In this respect, Asymgenesis provides a flexible charge-transfer framework with respect to the nature of the primordial asymmetry. The right part of Fig. 1 presents a schematic summary of charge transfer in Asymgenesis, with further discussion on its relation to other ADM frameworks provided in Appendix A. Finally, it is important to note that Asymgenesis serves solely as a framework for charge transfer between the visible and dark sector(s) in ADM models. Accordingly, any model implementation of Asymgenesis must comply with the usual phenomenological constraints on ADM, including those arising from the annihilation of symmetric dark matter components and from washout processes induced by particle–antiparticle oscillations in the dark sector.
@@
[…]
Figure 2: Schematic evolution of chemical potentials in Asymgenesis for $T_{\rm{eq}}>T_{B-L}$ (left) and $T_{B-L}>T_{\rm{eq}}$ (right). Here, $T_{\rm{CG}},\,T_{B-L},\,T_{\rm{eq}},\,T_{C},$ and $T_{\rm{EW}}$ denote the temperature scales associated with chargegenesis, RHN interactions becoming inefficient, decoupling of charge transfer between the dark and visible sectors, $C$-violating interactions becoming efficient, and the electroweak phase transition, respectively.
Warmup —  We initiate our analysis by examining the transfer of charge between the visible and the dark sector. The charge asymmetry $q_{i}$ of a SM or DM particle species $i$ is defined as, $q_{i}=n_{i}-n_{\bar{i}}=g_{i}\mu_{i}T^{2}/6$, [^3] This formula is valid in the limit $T\gg m_{i}$, and is justified by the fact that all relevant charge-generation and charge-transfer processes must occur at $T\gtrsim 100$ TeV in the scenario to be considered and that $m_{X}\lesssim 10$ TeV from perturbative unitarity. where $n_{i}$ $(n_{\bar{i}})$ denotes the (anti)particle number density of particle $i$ with chemical potential $\mu_{i}$ and multiplicity $g_{i}$. At high temperatures, the interaction described in Eq. (1) may establish thermal equilibrium between the two sectors, relating their chemical potentials $\mu_{i}$ Harvey and Turner (1990). Further, all SM and DM chemical potentials can be expressed as linear combinations of a set of linearly independent conserved charges $q_{C}=\bar{\mu}_{C}T^{2}/6$, e.g. by employing the linear-algebra formalism in Domcke et al. (2021) (see also Antaramian et al. (1994); Fong (2016) for some related earlier works). Here, $\bar{\mu}_{C}$ is introduced for notional convenience: $\bar{\mu}_{C}$ and $q_{C}$ are not conjugate to each other in the usual thermodynamical sense, but both describe the conserved charge. In other words, $\bar{\mu}_{C}$ corresponds to $q_{C}$, written in units of a chemical potential. To illustrate the concept of Asymgenesis, we consider a specific example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$, with $L_{e}$, $H$, and $e_{R}$ representing the first-generation lepton doublet, the SM Higgs doublet, and the right-handed electron, respectively. If the higher-dimensional operator is in thermal equilibrium, the following equilibrium condition relating the SM and the dark sector holds,
@@
While the operator $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ violates right-handed electron number, the interaction Eq. (1) conserves a combination of right-handed electron number and dark number. Consequently, the conservation of the right-handed electron charge, $q_{e}\sim\mu_{e}/T$, at high temperatures Campbell et al. (1992); Cline et al. (1993, 1994); Bödeker and Schröder (2019) is modified as follows
@@
where we used $g_{X}=2$, which we employ throughout this paper. As an example, if the charge-transfer interaction decouples at temperatures of a few $100$ TeV, then the ADM chemical potential can be written as
@@
The ADM abundance would in this case be determined by two factors: (i) the primordial asymmetries in $q_{e}$ and $q_{X}$, and (ii) the values of the flavored BLB-L asymmetries, $\Delta_{\alpha}$, at the time of charge-transfer decoupling. If the lightest RHN, $N_{1}$, has a mass around $M_{N}\sim 10^{2-3}$ TeV and decouples while the portal interaction is still efficient, then WILG in the strong wash-in regime would lead to